How do I write the function for 850 x3.5% increase each year?

Answer by Alan3354(69443)

How do I write the function for 850 x3.5% increase each year?
------------
FV = future value
PV = present value = 850
3.5% = 0.035
---
FV = PV*(1 = 0.035)^t, t = # of years
FV = PV*1.035^t